# TTS-VAR: A Test-Time Scaling Framework for Visual Auto-Regressive Generation

### v1

@@ Abstract
Scaling visual generation models is essential for real-world content creation, yet requires substantial training and computational expenses. Alternatively, test-time scaling has garnered growing attention due to resource efficiency and promising performance. In this work, we present TTS-VAR, the first general test-time scaling framework for visual auto-regressive (VAR) models, modeling the generation process as a path searching problem. To dynamically balance computational efficiency with exploration capacity, we first introduce an adaptive descending batch size schedule throughout the causal generation process. Besides, inspired by VAR’s hierarchical coarse-to-fine multi-scale generation, our framework integrates two key components: (i) At coarse scales, we observe that generated tokens are hard for evaluation, possibly leading to erroneous acceptance of inferior samples or rejection of superior samples. Noticing that the coarse scales contain sufficient structural information, we propose clustering-based diversity search. It preserves structural variety through semantic feature clustering, enabling later selection on samples with higher potential. (ii) In fine scales, resampling-based potential selection prioritizes promising candidates using potential scores, which are defined as reward functions incorporating multi-scale generation history. Experiments on the powerful VAR model Infinity show a notable 8.7% GenEval score improvement (0.69→0.75). Key insights reveal that early-stage structural features effectively influence final quality, and resampling efficacy varies across generation scales. Code is available at https://github.com/ali-vilab/TTS-VAR.
@@ 1 Introduction
Inspired by this perspective, we explore whether VAR models can also benefit from path searching. However, directly adopting the two strategies from diffusion models is non-trivial. For the former, the extra inference steps incur prohibitively high computation for VAR models with exponential complexity growth. They also disrupt the KV Cache mechanism [28, 29], which is crucial for retaining efficiency in AR inference. The latter also fails to reach the expectation. We observe that the rewards of images at early scales struggle to accurately represent the quality of final images, leading to incorrectly ruling out certain early-scale tokens which could be promising for later-scale generation, as demonstrated in Sec. 5.3. We attribute this to the difference between VAR and diffusion models. In VAR, unlike diffusion process that can refine generated noise through iterative denoising, all tokens remain fixed once generated. Each token not only contributes to decoding the final image but also directly affects all subsequent token generation, resulting in much lower tolerance for poor early-stage tokens.
In this paper, we introduce the first Test-Time Scaling framework for VAR, abbreviated as TTS-VAR. Different from simple selection according to reward functions, we design scaling strategies aligned to the causal coarse-to-fine generation process of VAR. Firstly, noticing the progressively increased consumption of FLOPs and RAM in VAR, we implement our framework under an adaptive descending batch size schedule, reduced from larger batch sizes at coarse scales to smaller ones in fine scales. This promotes the expression of more possibilities with little additional consumption. Secondly, though early scales are hard to evaluate by reward functions, we observe that structural information, which has a great impact on image contents, can be captured since early scales, as shown in Fig. 1 (b) and Sec. 5.4. This motivates splitting the generation process into two key components: clustering-based diversity search for early scales and resampling-based potential selection for late scales. At early scales, while results of intermediate states can hardly be estimated, we aim to keep the diversity as batch size decreases, thus enabling later selections on samples with higher potential. We employ clustering on semantic features extracted by pre-trained extractors like DINOv2 [30], and pick from each category for dissimilar samples to ensure sampling diversity. At late scales, while scores of intermediate images share high consistency with those of final images, we calculate potential scores to directly resample preferred samples. The potential scores are specifically defined reward functions based on the generation history of all scales, instead of only the current one.
@@ 2.1 Test-time Scaling in Diffusion Models
Diffusion models [6, 31, 32, 33, 34, 35] create high-resolution images by denoising a Gaussian distribution into an image distribution. Initial research efforts [36, 5, 37] concentrated on scaling up the number of denoising steps to enhance the quality. However, it has been observed that as the number of inference steps increased, performance plateaued, and sampling additional steps is ineffective. Consequently, early research [38, 39, 40] mainly aims to reduce inference steps while maintaining image quality.
@@ 2.2 Test-time Scaling in Autoregressive Models
In autoregressive Large Language Models [9, 7, 8], test-time scaling is a widely employed technique to enhance performance. Since Wei et al. [15] proposed Chain-of-Thought and enabled LLMs to benefit from a structured thinking process, various studies [16, 17, 18, 19, 41] have explored tree search, graph search, and other methodologies to improve outcomes further. All these strategies leverage the reasoning capabilities of models and utilize the properties inherent in natural language for scaling.
However, within autoregressive image generative models, characterized by a deterministic process with a steady token length, it is unnatural to directly increase the image token sequence as "thinking". Instead, Guo et al. [20] conceptualize generation CoT as an image-level problem. By employing a unified understanding and generation model [42] that first generates and then evaluates, it self-corrects results to align with expectations. Nevertheless, this approach relies solely on evaluating results, neglecting the generation process itself. Jiang et al. [21], instead, propose splitting the task into semantic-level and token-level phases, enabling a multi-stage generation as the thinking process. However, this method necessitates additional reinforcement learning for fine-tuning.
@@ 3 Preliminary: Visual Auto-Regressive Modeling
Unlike traditional next-token prediction auto-regressive models such as LLama-Gen [10], Visual Auto-Regressive Modeling (VAR) [13] tokenizes an input image $I$ into a feature map $F\in\mathbb{R}^{h\times w\times d}$. With quantizer $Q$, it quantizes the feature map $F$ into a sequence of multi-scale discrete residual feature maps [43] $\{r_{i}\}_{i=1}^{K}$, where $K$ represents the number of residual features across varying resolutions. For each residual feature map $r_{k}$, the resolution is $h_{k}\times w_{k}$, which progressively increases from $k=1$ to $k=K$. Specifically, when $k=1$, $h_{k}=w_{k}=1$, and when $k=K$, $h_{k}=h,w_{k}=w$. From the sequence of residual features, at each scale $k$, a gradually refined feature map $f_{k}$ can be computed as:
$f_{k}=\sum_{i=1}^{k}\text{up}(r_{i},(h,w)),$ (1)
where $\text{up}(\cdot,\cdot)$ denotes upsampling the single-scale feature map to the target resolution, and $f_{k}$ is the aggregated sum of features $\{r_{i}\}_{i=1}^{k}$. During inference, the downsampled accumulated feature map $\tilde{f}_{k}=\text{down}(f_{k},(h_{k+1},w_{k+1}))$ is appended as initial tokens for the prediction of the next scale and a scale-wise causal mask is employed to facilitate local bi-directional information modeling. The transformer is trained to predict the next-scale residual feature map. In Infinity [14], a VAR-based model scaled up for text-to-image generation, the quantizer $Q$ is advanced from VQ [44] to BSQ [45]. Additionally, a Flan-T5 [46] text encoder $\Psi$ is harnessed for prompt embeddings. With text prompt $c$ as the condition, the overall likelihood is:
@@ 4 Method
In TTS-VAR, we conceptualize the generation of high-quality and human-preferred images as a path searching problem, and identify two primary subproblems: (i) how to search for more possibilities, and (ii) how to select intermediate states for superior final results. Besides applying an adaptive batch size schedule as illustrated in Sec. 4.1 to enlarge the search scope, we introduce clustering-based diversity search in Sec. 4.2 and resampling-based potential selection in Sec. 4.3 to solve these problems. The complete method is illustrated in Fig. 1 (c).
@@ 4.1 Adaptive Batch Sampling
Influenced by the causal attention mechanism, in VAR models, both RAM memory consumption and computational expense increase along with the length of the token sequence. As illustrated by the blue lines in Fig. 2 (a) and (b), during the inference process of early scales, memory requirements and computational costs are minimal. However, at later scales, when the pre-existing sequence extends significantly and the current prediction scale encompasses a large number of tokens, the resource consumption becomes substantial.
[…]
@@ 4.2 Clustering-Based Diversity Search
As sequence length increases, maintaining a large batch size becomes cost-prohibitive, necessitating a filtering method for the desired samples. A straightforward approach involves calculating the reward function for intermediate results and selecting those with higher scores. However, our findings in Sec. 5.3 indicate that rating models struggle to evaluate early intermediate images for reward scores consistent with the final images, as also observed by Guo et al. [20]. To avoid erroneous elimination of certain initial samples that may hold potential for later-scale generation, we explore ways to keep the diversity of samples.
We observe that during the generation process, unlike the details appearing in late scales, the structural information, which significantly influences the quality of final images, is conveyed from the early phases. We analyze this phenomenon in Sec. 5.4 and find extractors like DINOv2 [30] can effectively capture features strongly connected with structures. Given this, we create clusters based on the semantic features to filter samples from each category and ensure structural diversity, thereby maximally enhancing possibilities for valuable results.
Specifically, from the current batch size $b_{i}$, we need to select $b_{i+1}$ samples as next states. Firstly, for $b_{i}$ intermediate images $\{I_{j}\}_{j=1}^{b_{i}}$, each image is embedded in a high-dimensional semantic embedding space via a feature extractor $F$, creating a set of embeddings $S=\bigcup_{b_{i}}F(I_{j})$. Subsequently, we apply the K-Means++ [47] algorithm to cluster these embeddings into $b_{i+1}$ cluster centers and select samples with the shortest L2 distance to cluster centers as new batches.
To extract structural information for diversity, we primarily employ the self-supervised DINOv2 [30] as the extractor, generating the feature map $s\in\mathbb{R}^{(h^{\prime}\times w^{\prime})\times d}$. To obtain one-dimensional features for clustering, we apply PCA reduction on feature patches for $s^{\prime}\in\mathbb{R}^{(h^{\prime}\times w^{\prime})}$. We also consider pooling the second dimension for $s^{\prime}\in\mathbb{R}^{d}$, and supervised features from InceptionV3 [48] without the final fully connected layer. We discuss these choices in Sec. 5.4.
@@ 4.3 Resampling-Based Potential Selection
Different potential scores benefit distinct attributes of the generation history. For instance, DIFF favors samples with higher growth rates, while MAX favors those with higher ceilings. In our setting, VALUE performs well. We will explore these choices in Sec. 5.3.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of our TTS-VAR on the powerful VAR model Infinity-2B [14] with resampling temperature $\lambda=10$. We present the comparisons in Sec. 5.1 and Sec. 5.2, and precisely analyze design details in Sec. 5.3 and Sec. 5.4. Following previous work [24, 22, 52], we utilize ImageReward [53] as the reward function for guidance. We evaluate results using the main metric GenEval [54] and T2I-CompBench [55], with relevant indicators ImageReward [53], HPSv2.1 [56, 57], Aesthetic V2.5 [58], and CLIP-Score [59, 60], based on prompts offered by GenEval.
@@ 5.1 Overall Performance
Figure 3: Score Curves over Sample Number $N$.
@@
Infinity [14] 0.5688 0.7421 0.4557 0.6034 0.2279 0.4023 0.5479 0.7820 0.7890
Infinity+IS ($N=8$) 0.5965 0.7746 0.5078 0.6501 0.2462 0.4194 0.6002 0.7803 0.7937
Infinity+BoN ($N=8$) 0.6115 0.7950 0.5439 0.6886 0.2545 0.4205 0.6090 0.7870 0.7937
Infinity+Ours ($N=2$) 0.6151 0.7887 0.5578 0.6858 0.2697 0.4286 0.6112 0.7853 0.7936
Infinity+Ours ($N=8$) 0.6230 0.8073 0.5914 0.7121 0.2644 0.4302 0.6340 0.7880 0.7963
@@
We further evaluated performance on T2I-CompBench [55]. As presented, TTS-VAR exhibits a remarkable improvement across every indicator compared to the base model. Consistent with the consequences on GenEval, our method achieves superior results at $N=2$ compared to Best-of-N at $N=8$, and secures the highest scores on most individual items and the overall average.
@@ 5.2 Qualitative Comparison
We present images generated by different methods here for reference, to further explain the improvement in image quality and text alignment. As displayed in Fig. 4, our method correctly generated objects with the required number, for example, the "kayaks" in the first case and the "birds" in the second case, which is hard for base models and other scaling strategies to achieve. In the complex scenes with several pairs of numbers and colors, TTS-VAR ensures the color attribute like the "golden apples" in the fourth case, avoiding the problem of attribute forgetting.

### v2

@@ Abstract
Scaling visual generation models is essential for real-world content creation, yet requires substantial training and computational expenses. Alternatively, test-time scaling has garnered growing attention due to resource efficiency and promising performance. In this work, we present TTS-VAR, the first general test-time scaling framework for visual auto-regressive (VAR) models, modeling the generation process as a path searching problem. To dynamically balance computational efficiency with exploration capacity, we first introduce an adaptive descending batch size schedule throughout the causal generation process. Besides, inspired by VAR’s hierarchical coarse-to-fine multi-scale generation, our framework integrates two key components: (i) At coarse scales, we observe that generated tokens are hard for evaluation, possibly leading to erroneous acceptance of inferior samples or rejection of superior samples. Noticing that the coarse scales contain sufficient structural information, we propose clustering-based diversity search. It preserves structural variety through semantic feature clustering, enabling later selection on samples with higher potential. (ii) In fine scales, resampling-based potential selection prioritizes promising candidates using potential scores, which are defined as reward functions incorporating multi-scale generation history. Experiments on the powerful VAR model Infinity2B show a notable 8.7% GenEval score improvement (0.69→0.75). Key insights reveal that early-stage structural features effectively influence final quality, and resampling efficacy varies across generation scales. Code is available at https://github.com/ali-vilab/TTS-VAR.
@@ 1 Introduction
Inspired by this perspective, we explore whether VAR models can also benefit from path searching. However, directly adopting the two strategies from diffusion models is non-trivial. For the former, the extra inference steps incur prohibitively high computation for VAR models with exponential complexity growth. They also disrupt the KV Cache mechanism [28, 29], which is crucial for retaining efficiency in AR inference. The latter also fails to reach the expectation. We observe that the rewards of images at early scales struggle to accurately represent the quality of final images, leading to incorrectly ruling out certain early-scale tokens which could be promising for later-scale generation, as demonstrated in Sec.˜5.4. We attribute this to the difference between VAR and diffusion. In VAR, unlike diffusion process that can refine generated noise through iterative denoising, all tokens remain fixed once generated. Each token not only contributes to decoding the final image but also directly affects all subsequent token generation, resulting in much lower tolerance for poor early-stage tokens.
In this paper, we introduce the first Test-Time Scaling framework for VAR, abbreviated as TTS-VAR. Different from simple selection according to reward functions, we design scaling strategies aligned to the causal coarse-to-fine generation process of VAR. Firstly, noticing the progressively increased consumption of FLOPs and RAM in VAR, we implement our framework under an adaptive descending batch size schedule, reduced from larger batch sizes at coarse scales to smaller ones in fine scales. This promotes the expression of more possibilities with little additional consumption. Secondly, though early scales are hard to evaluate by reward functions, we observe that structural information, which has a great impact on image contents, can be captured since early scales, as shown in Fig.˜1 (b) and Sec.˜5.5. This motivates splitting the generation process into two key components: clustering-based diversity search for early scales and resampling-based potential selection for late scales. At early scales, while results of intermediate states can hardly be estimated, we aim to keep the diversity as batch size decreases, thus enabling later selections on samples with higher potential. We employ clustering on semantic features extracted by pre-trained extractors like DINOv2 [30], and pick from each category for dissimilar samples to ensure sampling diversity. At late scales, while scores of intermediate images share high consistency with those of final images, we calculate potential scores to directly resample preferred samples. The potential scores are specifically defined reward functions based on the generation history of all scales, instead of only the current one.
@@ 2.1 Test-time Scaling in Diffusion Models
Diffusion models [6, 31, 32, 33, 34] create high-resolution images by denoising a Gaussian distribution into an image distribution. Initial research efforts [35, 5, 36] concentrated on scaling up the number of denoising steps to enhance the quality. However, it has been observed that as the number of inference steps increased, performance plateaued, and sampling additional steps is ineffective. Consequently, early research [37, 38, 39] mainly aims to reduce inference steps while maintaining image quality.
@@ 2.2 Test-time Scaling in Autoregressive Models
In autoregressive Large Language Models [9, 7, 8], test-time scaling is a widely employed technique to enhance performance. Since Wei et al. [15] proposed Chain-of-Thought and enabled LLMs to benefit from a structured thinking process, various studies [16, 17, 18, 19] have explored tree search, graph search, and other methodologies to improve outcomes further. All these strategies leverage the reasoning capabilities of models and utilize the properties inherent in natural language for scaling.
However, within autoregressive image generative models [10, 13, 11, 40], characterized by a deterministic process with a steady token length, it is unnatural to directly increase the image token sequence as "thinking". Instead, Guo et al. [20] conceptualize generation CoT as an image-level problem. By employing a unified understanding and generation model [41] that first generates and then evaluates, it self-corrects results to align with expectations. Nevertheless, this approach relies solely on evaluating results, neglecting the generation process itself. Jiang et al. [21], instead, propose splitting the task into semantic-level and token-level phases, enabling a multi-stage generation as the thinking process. However, this method necessitates additional reinforcement learning for fine-tuning.
@@ 3 Preliminary: Visual Auto-Regressive Modeling
Unlike traditional next-token prediction auto-regressive models such as LLama-Gen [10], Visual Auto-Regressive Modeling (VAR) [13] tokenizes an input image $I$ into a feature map $F\in\mathbb{R}^{h\times w\times d}$. With quantizer $Q$, it quantizes the feature map $F$ into a sequence of multi-scale discrete residual feature maps [42] $\{r_{k}\}_{k=1}^{K}$, where $K$ represents the number of residual features across varying resolutions. For each residual feature map $r_{k}$, the resolution is $h_{k}\times w_{k}$, which progressively increases from $k=1$ to $k=K$. Specifically, when $k=1$, $h_{k}=w_{k}=1$, and when $k=K$, $h_{k}=h,w_{k}=w$. From the sequence of residual features, at each scale $k$, a gradually refined feature map $f_{k}$ can be computed as:
$f_{k}=\sum_{i=1}^{k}\text{up}(r_{k},(h,w)),$ (1)
where $\text{up}(\cdot,\cdot)$ denotes upsampling the single-scale feature map to the target resolution, and $f_{k}$ is the aggregated sum of features $\{r_{k}\}_{k=1}^{K}$. During inference, the downsampled accumulated feature map $\tilde{f}_{k}=\text{down}(f_{k},(h_{k+1},w_{k+1}))$ is appended as initial tokens for the prediction of the next scale and a scale-wise causal mask is employed to facilitate local bi-directional information modeling. The transformer is trained to predict the next-scale residual feature map. In Infinity [14], a VAR-based model scaled up for text-to-image generation, the quantizer $Q$ is advanced from VQ [43] to BSQ [44]. Additionally, a Flan-T5 [45] text encoder $\Psi$ is harnessed for prompt embeddings. With text prompt $c$ as the condition, the overall likelihood is:
@@ 4 Method
In TTS-VAR, we conceptualize the generation of high-quality and human-preferred images as a path searching problem, and identify two primary subproblems: (i) how to search for more possibilities, and (ii) how to select intermediate states for superior final results. Besides applying an adaptive batch size schedule as illustrated in Sec.˜4.1 to enlarge the search scope, we introduce clustering-based diversity search in Sec.˜4.2 and resampling-based potential selection in Sec.˜4.3 to solve these problems. The complete method is illustrated in Fig.˜1 (c).
@@ 4.1 Adaptive Batch Sampling
Influenced by the causal attention mechanism, in VAR models, both RAM memory consumption and computational expense increase along with the length of the token sequence. As illustrated by the blue lines in Fig.˜2 (a) and (b), during the inference process of early scales, memory requirements and computational costs are minimal. However, at later scales, when the pre-existing sequence extends significantly and the current prediction scale encompasses a large number of tokens, the resource consumption becomes substantial.
Therefore, we implement adaptive batch sizes during inference, capitalizing on the efficiency of lower consumption at early scales. This descending adaptive batch size schedule $\{b_{0},b_{1},\dots,b_{K}\}$, where $K$ represents the number of scales, generates more samples in earlier stages and fewer samples in the later stages. For a typical VAR model encompassing 13 scales, the batch size schedule is {8NN, 8NN, 6NN, 6NN, 6NN, 4NN, 2NN, 2NN, 2NN, 1NN, 1NN, 1NN, 1NN} unless otherwise specified. At early scales, clustering in Sec.˜4.2 filters several categories. At late scales, resampling in Sec.˜4.3 chooses superior states. As demonstrated in Fig.˜2, while the increased number of batches amplifies memory and computation costs at early scales, these additions are relatively minor in the overall expenditure.
@@ 4.2 Clustering-Based Diversity Search
As sequence length increases, maintaining a large batch size becomes cost-prohibitive, necessitating a filtering method for the desired samples. A straightforward approach involves calculating the reward function for intermediate results and selecting those with higher scores. However, our findings in Sec.˜5.4 indicate that rating models struggle to evaluate early intermediate images for reward scores consistent with the final images, as also observed by Guo et al. [20]. To avoid erroneous elimination of certain initial samples that may hold potential for later-scale generation, we explore ways to keep the diversity of samples.
We observe that during the generation process, unlike the details appearing in late scales, the structural information, which significantly influences the quality of final images, is conveyed from the early phases. We analyze this phenomenon in Sec.˜5.5 and find extractors like DINOv2 [30] can effectively capture features strongly connected with structures. Given this, we create clusters based on the semantic features to filter samples from each category and ensure structural diversity, thereby maximally enhancing possibilities for valuable results.
Specifically, from the current batch size $b_{i}$, we need to select $b_{i+1}$ samples as next states. Firstly, for $b_{i}$ intermediate images $\{I_{j}\}_{j=1}^{b_{i}}$, each image is embedded in a high-dimensional semantic embedding space via a feature extractor $F$, creating a set of embeddings $S=\bigcup_{j}F(I_{j})$. Subsequently, we apply the K-Means++ [46] algorithm to cluster these embeddings into $b_{i+1}$ cluster centers and select samples with the shortest L2 distance to cluster centers as new batches.
To extract structural information for diversity, we primarily employ the self-supervised DINOv2 [30] as the extractor, generating the feature map $s\in\mathbb{R}^{(h^{\prime}\times w^{\prime})\times d}$. To obtain one-dimensional features for clustering, we apply PCA reduction on feature patches for $s^{\prime}\in\mathbb{R}^{(h^{\prime}\times w^{\prime})}$. We also consider pooling the second dimension for $s^{\prime}\in\mathbb{R}^{d}$, and supervised features from InceptionV3 [47] without the final fully connected layer. We discuss these choices in Sec.˜5.5.
@@ 4.3 Resampling-Based Potential Selection
Different potential scores benefit distinct attributes of the generation history. For instance, DIFF favors samples with higher growth rates, while MAX favors those with higher ceilings. In our setting, VALUE performs well. We will explore these choices in Sec.˜5.4.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of our TTS-VAR on the powerful VAR model Infinity [14] with resampling temperature $\lambda=10$. We present the comparisons in Sec.˜5.1 and Sec.˜5.2, and precisely analyze design details in Sec.˜5.4 and Sec.˜5.5. Following previous work [24, 22, 51], we utilize ImageReward [52] as the reward function for guidance. We evaluate results using the main metric GenEval [53], and T2I-CompBench [54], with relevant indicators ImageReward [52], HPSv2.1 [55, 56], Aesthetic V2.5 [57], and CLIP-Score [58, 59], based on prompts offered by GenEval.
@@ 5.1 Overall Performance
Infinity 8B 0.8866 0.7292 0.6750 0.7646
Infinity+BoN ($N=4$) 8B 0.9175 0.7946 0.7225 0.7995
Infinity+Ours ($N=2$) 8B 0.9330 0.7798 0.7050 0.7985
Infinity+Ours ($N=4$) 8B 0.9304 0.8036 0.7600 0.8188
@@
Figure 3: Score Curves over Sample Number NN.
@@
To better demonstrate the generalizability of our framework across different models, we also conducted comparative experiments on the Infinity8B model. The results show that even on this stronger 8B model, our approach effectively boosts performance, improving accuracy from 0.76 to 0.82—an increase of about 7Moreover, it surpasses the gains achieved by methods such as IS and BoN under the same settings. This indicates that our method can be directly applied to other VAR-based architectures.
@@
Infinity2B [14] 0.5688 0.7421 0.4557 0.6034 0.2279 0.4023 0.5479 0.7820 0.7890
Infinity2B+IS ($N=8$) 0.5965 0.7746 0.5078 0.6501 0.2462 0.4194 0.6002 0.7803 0.7937
Infinity2B+BoN ($N=8$) 0.6115 0.7950 0.5439 0.6886 0.2545 0.4205 0.6090 0.7870 0.7937
Infinity2B+Ours ($N=2$) 0.6151 0.7887 0.5578 0.6858 0.2697 0.4286 0.6112 0.7853 0.7936
Infinity2B+Ours ($N=8$) 0.6230 0.8073 0.5914 0.7121 0.2644 0.4302 0.6340 0.7880 0.7963
@@
We further evaluated performance on T2I-CompBench [54]. As presented in Table˜2, TTS-VAR exhibits a remarkable improvement across every indicator compared to the base model. Consistent with the consequences on GenEval, our method achieves superior results at $N=2$ compared to Best-of-N at $N=8$, and secures the highest scores on most individual items and the overall average.
@@ 5.2 Qualitative Comparison
We present images generated by different methods here for reference, to further explain the improvement in image quality and text alignment. As displayed in Fig.˜4, our method correctly generated objects with the required number, for example, the "kayaks" in the first case and the "birds" in the second case, which is hard for base models and other scaling strategies to achieve. In the complex scenes with several pairs of numbers and colors, TTS-VAR ensures the color attribute like the "golden apples" in the fourth case, avoiding the problem of attribute forgetting.
5.3 User Study
To evaluate from the view of practical usage and gather user feedback, we selected a small set of 15 image groups and compiled them into a questionnaire on Google Forms. For each group, the images are randomly shuffled, and participants were asked to select the best generated image based on three criteria: Image Quality, Image Rationality, and Consistency with the Prompt. In total, we collected 21 completed questionnaires, resulting in 315 sets of survey data for each indicator. The percentages of votes received by each approach are summarized in Table˜3 below.
Metric Baseline IS BoN Ours
Image Quality 13.3% 7.9% 13.3% 65.4%
Image Rationality 13.7% 8.6% 8.6% 69.2%
Consistency with the Prompt 1.3% 1.9% 2.5% 94.3%
Table 3: User Study.